OpenStudy (anonymous):

(-y/2z)(-6z/y) simplify plz help i keep getting 3!

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

the minus signs go the y 's cancel the z's cancel and two goes in to six three times

OpenStudy (anonymous):

well my teacher marked it wrong so i thought that i was wrong

OpenStudy (anonymous):

\[\frac{ -y }{ 2z }\frac{ -6z }{ y }=\frac{ 6z }{ 2z }=3\]

myininaya (myininaya):

\[\frac{-y}{2z} \frac{-6z}{y}=\frac{y}{2z}\frac{6z}{y}=\frac{6}{2}=3\] when y and z aren't zero --- also did you write 3! or 3?

OpenStudy (anonymous):

3